OpenStudy (anonymous):

solve for x: 3/x - 2/3 = 5/6

OpenStudy (anonymous):

x = 3?

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

\[\frac{ 3 }{ x }=\frac{ 5 }{ 6 }+\frac{ 2 }{ 3 }\] \[\frac{ 3 }{x }=\frac{ 9 }{ 6 }\] \[\frac{ 1 }{ x }=\frac{ 1 }{ 2 }\] x=2